OpenStudy (anonymous):

How do you graph x^2- 3y^2=-11?

OpenStudy (dumbcow):

get into standard form of a hyperbola \[\frac{(x-h)^{2}}{a^{2}}-\frac{(y-k)^{2}}{b^{2}} = 1\]

OpenStudy (anonymous):

Like a hyperbola.

OpenStudy (dumbcow):

Divide by -11 \[ -\frac{x^{2}}{11} -(\frac{-3}{11})y^{2} = 1 \rightarrow \frac{y^{2}}{\frac{11}{3}} - \frac{x^{2}}{11} =1\]

OpenStudy (dumbcow):

center: (0,0) \[a^{2} = \frac{11}{3}\rightarrow a = \sqrt{\frac{11}{3}}\] \[b^{2} = 11 \rightarrow b = \sqrt{11}\] vertices: (0, +-a) \[c^{2} = a^{2} +b^{2} = \frac{11}{3}+11 = \frac{44}{3}\] \[c = \sqrt{\frac{44}{3}}\] foci: (0, +-c)

OpenStudy (dumbcow):

were you able to finish the graph?................ do you have to do the slant asymptotes also?